# Leptogenesis with perturbations in type-II and type-III seesaw models

### v1

@@ Abstract
Density perturbations have recently been shown to lead to a novel effect in the freeze-out of heavy particles called “acoustically driven freeze-out.” This leads to an enhancement in the yield in standard leptogenesis. We extend this calculation to include $2\to 2$ washout processes in type-I leptogenesis and the Sommerfeld-enhanced $2\to 2$ gauge annihilations in type-II and type-III leptogenesis. These CP conserving annihilations suppress the yield of heavy particles sourcing the asymmetry. We show the acoustically driven freeze-out leads to novel enhancements in the baryon asymmetry in type-II and type-III leptogenesis, already in the weak washout regime, in contrast with type-I leptogenesis.
@@ I Introduction
One objection to this whole enterprise may be that without a reason for expecting significant enhancements in $\mathcal{R}_{i}$ — at length scales of order of the hubble horizon $\sim 1/H$ during the freeze-out process — these calculations are only of academic interest, and will be unlikely to be of relevance for explaining the observed universe (unlike the perturbations at CMB scales Senatore et al. (2009)). We note, however, that strong phase transitions associated with the generation of heavy DM Hambye and Strumia (2013); Hambye et al. (2018); Baldes and Garcia-Cely (2019); Baldes et al. (2021a, 2022); Kierkla et al. (2023); Wong and Xie (2023); Balan et al. (2025) or heavy (seesaw) states associated with baryogensis Lazarides et al. (1986); Buchmuller et al. (2012); Buchmüller et al. (2013); Brivio and Trott (2017, 2019); Brdar et al. (2019a, b, c); Brivio et al. (2019); Brdar et al. (2019d); Baldes et al. (2021b); Huang and Xie (2022); Dasgupta et al. (2022); Chun et al. (2023), will also lead to the generation of curvature perturbations with length scales $\sim 1/H$ Liu et al. (2023); Elor et al. (2024); Lewicki et al. (2024); Zou et al. (2025). If the reheating temperature is subsequently high enough for the DM or heavy seesaw states to come back into equilibrium, the generated perturbations may well then lead to acoustically driven freeze-out playing a role. This motivates the current study in which we further explore acoustically driven freeze-out in isolation from other effects. The calculation in a more complete scenario can then benefit from the understanding gained. Similarly, the role of all these phase transitions, inhomogeneities, and density perturbations in generating detectable gravitational waves as a further signal of such a scenario is left for future work.
@@ II Leptogenesis models
In all three of the above seesaw realizations, a $\mathcal{B}-\mathcal{L}$ asymmetry can be generated through CP violating decays of the heavy states, in a rather straight forward fashion. We consider the decays of the lightest of the heavy states, with mass $M_{1}$, and parameterize the total decay rate in the standard way. We write the effective contribution of this state to the neutrino mass as
$\tilde{m_{1}}$ $=\frac{(y_{N}y_{N}^{\dagger})v^{2}}{M_{1}},\quad\mathrm{[Type-I]},$ (6)
$\tilde{m_{1}}$ $=\frac{(y_{{TL}}y_{{TH}}^{\dagger})v^{2}}{M_{1}},\quad\mathrm{[Type-II]},$ (7)
$\tilde{m_{1}}$ $=\frac{(y_{\Sigma}y_{\Sigma}^{\dagger})v^{2}}{M_{1}},\quad\mathrm{[Type-III]},$ (8)
@@
$\Gamma_{N}=\frac{\tilde{m_{1}}M_{1}}{8\pi v^{2}},\quad\mathrm{[Type-I]},$ (9)
$\Gamma_{\Sigma}=\frac{\tilde{m_{1}}M_{1}}{8\pi v^{2}},\quad\mathrm{[Type-III]}.$ (10)
@@
$\Gamma(T\to\bar{L}\bar{L})\equiv B_{L}\Gamma_{T}$ $=\mathrm{Tr}(y_{{TL}}y_{{TL}}^{\dagger})\frac{M_{T}}{16\pi}$ (11)
$\Gamma(T\to HH)\equiv B_{H}\Gamma_{T}$ $=(y_{{TH}}y_{{TH}}^{\dagger})\frac{M_{T}}{16\pi},$ (12)
where $B_{L}+B_{H}=1$ are the branching fractions. For simplicity, when we scan over parameter space, solving the Boltzmann equations below, we will be considering the limit $B_{L}=B_{H}=0.5$. Then we also take the single generation case for $y_{{TL}}$, and have
$\Gamma_{T}=\frac{\tilde{m_{1}}M_{1}}{8\pi v^{2}},\quad\mathrm{[Type-II,\;Simplified]}$ (13)
@@ III Density and temperature perturbations
The perturbations we consider are identical to those introduced by Hotokezaka, Jinno, and Takada Hotokezaka et al. (2025). We follow their definitions and take $T=\bar{T}+\delta T$, where $\bar{T}$ is the spatially averaged temperature of the universe, and define the fluctuation as $\delta_{T}\equiv\delta T/\bar{T}$. As a “time” variable in the Boltzmann equations it is useful to define
$\bar{z}\equiv\frac{M_{1}}{\bar{T}}$ (21)
@@
We work in Newtonian gauge of linearized cosmological perturbation theory, with negligible anisotropic stress, so that the two scalar perturbations of the metric are equal $\Phi=\Psi$. In radiation domination, for a given Fourier mode with wavenumber $k=|\mathbf{k}|$, one has Hotokezaka et al. (2025)
@@ IV.2 Type-II Boltzmann equation
Due to their gauge interactions, the scalar triplets can annihilate into SM fermions, Higgs Bosons, and gauge bosons $T\bar{T}\to L\bar{L},H\bar{H},AA,YY$. The overall annihilation rate $\gamma_{{TA}}(z_{T})$ can be found using the reduced cross sections reported in (Hambye et al., 2006, Eqs. (16a)-(16d)). Sommerfeld enhancement is taken into account by following Strumia (2009). The gauge couplings are found at the appropriate scale, $\sim T$, using the RGEs in Buttazzo et al. (2013). The off-shell washout rates can be found by making use of the reduced cross sections in (Hambye et al., 2006, Eqs. (17a) and (17b)), with appropriate modifications to make use of the RIS subtracted propagator, using the procedure of Giudice et al. (2004), so that $\hat{\sigma}_{Ts}\to\hat{\sigma}_{Ts}^{\rm sub}$. (We also correct some trivial typos by replacing $M_{T}\to m_{T}$ inside the square brackets in (Hambye et al., 2006, Eqs. (17a) and (17b)). Then we identify $m_{T}$ with $\tilde{m}_{1}$ in the limit limit $B_{L}=B_{H}=0.5$ in our notation for when we solve the equations. We also set the parameter $m_{H}=0.05\;\mathrm{eV}$ — it is the contribution to the neutrino mass mediated by heavier particles, not the SM Higgs mass. The $M_{T}^{2}$ factor outside the square brackets in (Hambye et al., 2006, Eqs. (17a) and (17b)) is of course simply $M_{1}^{2}$ in our notation.)
@@
$[1-\Psi(\bar{z})]s(z_{T})H(\bar{z})\bar{z}\frac{d\Delta_{H}}{d\bar{z}}$ $=\epsilon_{T}\gamma_{T}(z_{T})\left[\frac{\Sigma_{T}}{\Sigma_{T}^{\rm eq}(z_{T})}-1\right]-2\gamma_{T}^{\rm sub}(z_{T})\left[\frac{\Delta_{L}}{Y_{L}^{\rm eq}}+\frac{\Delta_{H}}{Y_{H}^{\rm eq}}\right]-2B_{H}\gamma_{T}(z_{T})\left[\frac{\Delta_{H}}{Y_{H}^{\rm eq}}-\frac{\Delta_{T}}{\Sigma_{T}^{\rm eq}(z_{T})}\right],$
@@
Figure 3: Contours showing the enhancement of the yield with perturbations for the three seesaw models (with $B_{L}=B_{H}=0.5$ for type-II). As well established in the previous literature, preferred regions (green) assuming hierarchical spectrum CP violation limits, Eqs. (18)-(20), are at somewhat higher $M_{1}$ in the type-II and type-III scenarios, due to the gauge interactions keeping the number densities closer to equilibrium. (We are assuming initial thermal $N$ abundance for the type-I green region, which affects the shape of this region at small $\tilde{m}_{1}$.) The new result shows that in the presence of perturbations, i.e. acoustically driven freeze-out, gauge interactions allow yield enhancements in the weak washout regime, $\tilde{m_{1}}<m_{\ast}$. This occurs in parameter ranges where the decay time approximately matches the freeze-out time of the gauge interactions. At large values of $M_{1}$ the yields become suppressed due to off-shell $2\to 2$ washout interactions. For very large $M_{1}$ the couplings become non-perturbative (red shaded regions). Observations of the CMB limit the post-inflationary reheat temperature to $T_{\rm RH}<6.6\times 10^{15}\;\mathrm{GeV}$ Akrami et al. (2020); Ade et al. (2021) which gives a similar consistency constraint on the scenario, $M_{1}<T_{\rm RH}$, as perturbativity.
@@ V Results
Finally in Fig. 3, we collect our results, by showing standard plots of the preferred regions for leptogenesis in the different seesaw models, together with contours showing the enhancement of the yield for acoustically driven freeze-out, in terms of the parameters $M_{1}$ and $\tilde{m_{1}}$. (Note we take $B_{H}=B_{L}=0.5$ for type-II only for simplicity. Away from this limit, the washout rate decreases, but so does the amount of CP violation, so that the baryon yield is suppressed. There are also some differences in the preferred green region from CP violation at large $M_{1}$ and $\tilde{m}_{1}$ for type-II, compared with the results of (Strumia, 2009, Fig. 2), we were not able to find the source of this discrepancy.)
@@ VI Conclusion
We have extended the study of acoustically driven freeze out to include $2\to 2$ interactions, by examining in some detail leptogenesis in the type-II and type-III seesaw models in the presence of perturbations, and contrasting our results with those of type-I leptogenesis. Modest enhancements of the yield, $\sim 50\%$, are possible with primordial curvature perturbations $\mathcal{R}_{i}\sim 0.3$, already in the weak washout regime in the scenarios with gauge interactions. (The crucial ingredient being annihilations involving the species freezeing out in the external states — similar effects could be expected in extensions of the type-I seesaw involving annihilations of $N_{1}$ into scalars Aristizabal Sierra et al. (2014).)
@@
It would be of interest to study acoustically driven freeze-out in the context of dark matter, to see what effect, if any, is present in the freeze-out of stable species. Preliminary investigations suggest that in the standard DM freezeout scenario, modest enhancements of the required DM annihilation cross section $\sim 10\%$ are necessary to match the relic abundance, for perturbations $\mathcal{R}_{i}\sim 0.3$ Baldes (2025). It would also be of interest to see whether models featuring inflation, or strong early universe phase transitions, can be found which link physics of the particles undergoing freeze-out, with the generation of large $\mathcal{R}_{i}$ on the length scales $\sim 1/H(\bar{T}=M)$, relevant for acoustically driven freeze-out.

### v2

@@ Abstract
Density perturbations have recently been shown to lead to a novel effect in the freeze-out of heavy particles called “acoustically driven freeze-out.” This leads to an enhancement in the yield in standard leptogenesis. We extend this calculation to include $2\to 2$ washout processes in type-I leptogenesis and the Sommerfeld-enhanced $2\to 2$ gauge annihilations in type-II and type-III leptogenesis. These CP conserving gauge annihilations suppress the yield of heavy particles sourcing the asymmetry. We show the acoustically driven freeze-out leads to novel enhancements in the baryon asymmetry in type-II and type-III leptogenesis, already in the weak washout regime, in contrast with type-I leptogenesis.
@@ I Introduction
One objection to this whole enterprise may be that without a reason for expecting significant enhancements in $\mathcal{R}_{i}$ — at length scales of order of the hubble horizon $\sim 1/H$ during the freeze-out process — these calculations are only of academic interest, and will be unlikely to be of relevance for explaining the observed universe, unlike the perturbations at CMB scales Senatore et al. (2009) (for historical studies of small scale inhomogeneties during BBN, see e.g. Terasawa and Sato (1989); Kurki-Suonio and Matzner (1989)). We note, however, that strong phase transitions associated with the generation of heavy DM Hambye and Strumia (2013); Hambye et al. (2018); Baldes and Garcia-Cely (2019); Baldes et al. (2021a, 2022); Kierkla et al. (2023); Wong and Xie (2023); Balan et al. (2025) or heavy (seesaw) states associated with baryogensis Lazarides et al. (1986); Buchmuller et al. (2012); Buchmüller et al. (2013); Brivio and Trott (2017, 2019); Brdar et al. (2019a, b, c); Brivio et al. (2019); Brdar et al. (2019d); Baldes et al. (2021b); Huang and Xie (2022); Dasgupta et al. (2022); Chun et al. (2023), will also lead to the generation of curvature perturbations with length scales $\sim 1/H$ Liu et al. (2023); Elor et al. (2024); Lewicki et al. (2024); Zou et al. (2025). If the reheating temperature is subsequently high enough for the DM or heavy seesaw states to come back into equilibrium, the generated perturbations may well then lead to acoustically driven freeze-out playing a role.
This motivates the current study in which we further explore acoustically driven freeze-out in isolation from other effects. The calculation in a more complete scenario can then benefit from the understanding gained. Similarly, the role of all these phase transitions, inhomogeneities, and density perturbations in generating detectable gravitational waves as a further signal of such a scenario is left for future work.
Finally, we note that the isocurvature perturbations on observable CMB scales — which are super-horizon sized when sourced from out-of-equilibrium effects in our leptogenesis scenario, or from dark-matter freeze-out/in — will be highly suppressed and so not constrain the scenario Weinberg (2004); Bellomo et al. (2023); Strumia (2023); Racco and Riotto (2023); Holst et al. (2024); Stebbins (2023).
@@ II Leptogenesis models
In all three of the above seesaw realizations, a $\mathcal{B}-\mathcal{L}$ asymmetry can be generated through CP violating decays of the heavy states, in a rather straight forward fashion. We consider the decays of the lightest of the heavy states, with mass $M_{1}$, and parameterize the total decay rate in the standard way. We write the effective neutrino mass as Buchmuller et al. (2005); Hambye et al. (2006)
$\tilde{m_{1}}$ $=\frac{(y_{N}y_{N}^{\dagger})_{11}v^{2}}{M_{1}},\quad\mathrm{[Type-I]},$ (6)
$\tilde{m_{1}}$ $=\sqrt{(y_{{TH}}y_{{TH}}^{\dagger})\mathrm{Tr}(y_{{TL}}y_{{TL}}^{\dagger})}\frac{v^{2}}{M_{1}},\quad\mathrm{[Type-II]},$ (7)
$\tilde{m_{1}}$ $=\frac{(y_{\Sigma}y_{\Sigma}^{\dagger})_{11}v^{2}}{M_{1}},\quad\mathrm{[Type-III]},$ (8)
@@
$\Gamma_{N}$ $=\frac{\tilde{m_{1}}M_{1}^{2}}{8\pi v^{2}},\quad\mathrm{[Type-I]},$ (9)
$\Gamma_{\Sigma}$ $=\frac{\tilde{m_{1}}M_{1}^{2}}{8\pi v^{2}},\quad\mathrm{[Type-III]}.$ (10)
@@
$\Gamma(T\to\bar{L}\bar{L})$ $\equiv B_{L}\Gamma_{T}=\mathrm{Tr}(y_{{TL}}y_{{TL}}^{\dagger})\frac{M_{1}}{16\pi}$ (11)
$\Gamma(T\to HH)$ $\equiv B_{H}\Gamma_{T}=(y_{{TH}}y_{{TH}}^{\dagger})\frac{M_{1}}{16\pi},$ (12)
where $B_{L}+B_{H}=1$ are the branching fractions. For simplicity, when we scan over parameter space, solving the Boltzmann equations below, we will be considering the limit $B_{L}=B_{H}=0.5$. Then
$\Gamma_{T}=\frac{\tilde{m_{1}}M_{1}^{2}}{8\pi v^{2}},\quad\mathrm{[Type-II,\;Simplified]}.$ (13)
@@ III Density and temperature perturbations
The perturbations we consider are identical to those used by Hotokezaka, Jinno, and Takada Hotokezaka et al. (2025) and consist of radiation density perturbations oscillating in a radiation dominated universe. We follow their definitions and take $T=\bar{T}+\delta T$, where $\bar{T}$ is the spatially averaged temperature of the universe, and define the fluctuation as $\delta_{T}\equiv\delta T/\bar{T}$. As a “time” variable in the Boltzmann equations it is useful to define
$\bar{z}\equiv\frac{M_{1}}{\bar{T}}.$ (21)
@@
We work in Newtonian gauge of linearized cosmological perturbation theory, with negligible anisotropic stress, so that the two scalar perturbations of the metric are equal $\Phi=\Psi$. In radiation domination, for a given Fourier mode with wavenumber $k=|\mathbf{k}|$, one has Peter and Uzan (2013); Hotokezaka et al. (2025)
@@ IV.2 Type-II Boltzmann equation
Due to their gauge interactions, the scalar triplets can annihilate into SM fermions, Higgs Bosons, and gauge bosons $T\bar{T}\to L\bar{L},H\bar{H},AA,YY$. The overall annihilation rate $\gamma_{{TA}}(z_{T})$ can be found using the reduced cross sections reported in (Hambye et al., 2006, Eqs. (16a)-(16d)). Sommerfeld enhancement is taken into account by following Strumia (2009). The gauge couplings are found at the appropriate scale, $\sim T$, using the RGEs in Buttazzo et al. (2013). The off-shell washout rates can be found by making use of the reduced cross sections in (Hambye et al., 2006, Eqs. (17a) and (17b)), with appropriate modifications to make use of the RIS subtracted propagator, using the procedure of Giudice et al. (2004), so that $\hat{\sigma}_{Ts}\to\hat{\sigma}_{Ts}^{\rm sub}$. (We also correct some trivial typos by replacing $M_{T}\to\tilde{m}_{T}$ inside the square brackets in (Hambye et al., 2006, Eqs. (17a) and (17b)). Then we identify $\tilde{m}_{T}$ with $\tilde{m}_{1}$, in our notation. We also set the parameter $m_{H}=0.05\;\mathrm{eV}$ — it is the contribution to the neutrino mass mediated by heavier particles, not the SM Higgs mass. The $M_{T}^{2}$ factor outside the square brackets in (Hambye et al., 2006, Eqs. (17a) and (17b)) is of course simply $M_{1}^{2}$ in our notation.)
@@
$[1-\Psi(\bar{z})]s(z_{T})H(\bar{z})\bar{z}\frac{d\Delta_{H}}{d\bar{z}}$ $=\epsilon_{T}\gamma_{T}(z_{T})\left[\frac{\Sigma_{T}}{\Sigma_{T}^{\rm eq}(z_{T})}-1\right]-2\gamma_{T}^{\rm sub}(z_{T})\left[\frac{\Delta_{L}}{Y_{L}^{\rm eq}}+\frac{\Delta_{H}}{Y_{H}^{\rm eq}}\right]-2B_{H}\gamma_{T}(z_{T})\left[\frac{\Delta_{H}}{Y_{H}^{\rm eq}}-\frac{\Delta_{T}}{\Sigma_{T}^{\rm eq}(z_{T})}\right].$
@@
Figure 3: Contours showing the enhancement of the yield with perturbations for the three seesaw models (with $B_{L}=B_{H}=0.5$ for type-II). As well established in the previous literature, preferred regions (green) assuming hierarchical spectrum CP violation limits, Eqs. (18)-(20), are at somewhat higher $M_{1}$ in the type-II and type-III scenarios, due to the gauge interactions keeping the number densities closer to equilibrium. (We are assuming initial thermal $N$ abundance for the type-I green region, which affects the shape of this region at small $\tilde{m}_{1}$.) The new result shows that in the presence of perturbations, i.e. acoustically driven freeze-out, gauge interactions allow yield enhancements in the weak washout regime, $\tilde{m_{1}}<m_{\ast}$. This occurs in parameter ranges where the decay time approximately matches the freeze-out time of the gauge interactions. At large values of $M_{1}$ the yields become suppressed due to off-shell $2\to 2$ washout interactions. For very large $M_{1}$ the couplings become non-perturbative (red shaded regions).
@@ V Results
Finally in Fig. 3, we collect our results, by showing standard plots of the preferred regions for leptogenesis in the different seesaw models, together with contours showing the enhancement of the yield for acoustically driven freeze-out, in terms of the parameters $M_{1}$ and $\tilde{m_{1}}$. Cosmological observations disfavour $\sum m_{\nu i}>0.12\;\mathrm{eV}$ Aghanim et al. (2020); Alam et al. (2021), so also $\tilde{m}_{1}>0.12\;\mathrm{eV}$ is disfavoured Buchmuller et al. (2003), but we extend the plots to higher values for completeness. At large values of $M_{1}$ the couplings required to explain the neutrino masses become non-perturbative. Observations of the CMB limit the post-inflationary reheat temperature to $T_{\rm RH}<6.6\times 10^{15}\;\mathrm{GeV}$ Akrami et al. (2020); Ade et al. (2021) which gives a similar consistency constraint on the scenario, $M_{1}<T_{\rm RH}$, as perturbativity.
Note we take $B_{H}=B_{L}=0.5$ for type-II only for simplicity. Away from this limit, the washout rate decreases, but so does the amount of CP violation, so that the baryon yield is suppressed. (There are also some differences in the preferred green region from CP violation at large $M_{1}$ and $\tilde{m}_{1}$ for type-II, compared with the results of (Strumia, 2009, Fig. 2), we were not able to find the source of this discrepancy, e.g. by allowing for larger $\sum m_{\nu i}^{2}$ in Eq. (19), the bound on $|\epsilon_{T}|$.)
@@ VI Conclusion
We have extended the study of acoustically driven freeze out to include $2\to 2$ interactions, by examining in some detail leptogenesis in the type-II and type-III seesaw models in the presence of perturbations, and contrasting our results with those of type-I leptogenesis. Modest enhancements of the yield, $\sim 40\%$, are possible with primordial curvature perturbations $\mathcal{R}_{i}\sim 0.3$, already in the weak washout regime in the scenarios with gauge interactions. (The crucial ingredient being annihilations involving the species freezeing out in the external states — similar effects could be expected in extensions of the type-I seesaw involving annihilations of $N_{1}$ into scalars Aristizabal Sierra et al. (2014).)
@@
It would be of interest to study acoustically driven freeze-out in the context of dark matter, to see what effect, if any, is present in the freeze-out of stable species. Preliminary investigations suggest that in the standard DM freeze-out scenario, modest enhancements of the required DM annihilation cross section $\sim 10\%$ are necessary to match the relic abundance, for perturbations $\mathcal{R}_{i}\sim 0.3$ Baldes (2025) (for effects of small scale perturbations during freeze-in, see (Stebbins, 2023, Fig. 1)). It would also be of interest to see whether models featuring inflation, or strong early universe phase transitions, can be found which link physics of the particles undergoing freeze-out, with the generation of large $\mathcal{R}_{i}$ on the length scales $\sim 1/H(\bar{T}=M)$, relevant for acoustically driven freeze-out.